chapter 24 part 4

1. Approximately what is the smallest detail (in nanometers) observable with a microscope that uses light of frequency 0.61×1015 Hz?

2. Lasers can be constructed that produce an extremely high intensity electromagnetic wave for a brief time—called pulsed lasers. They are used to ignite nuclear fusion, for example. Such a laser may produce an electromagnetic wave with a maximum electric field strength of 1.94×1011 V / m for a time of 0.97 ns. What is the intensity of the beam in 1019 W/m2?
Answer #1

1.

We know that

frequency = f = c/lambda

lambda = c/f

lambda = 3*10^8/(0.61*10^15) = 491.8*10^-9 m

lambda = 491.8 nm

2.

Intensity is given by:

I = c*e0*E0^2/2

Using given values:

I = 3*10^8*8.85*10^-12*(1.94*10^11)^2/2

I = 4.99*10^19 W/m^2

Please Upvote.